OpenStudy (anonymous):

what is the 12 term in the sequence -18/5,6/5,-2/5,2/15 an=-18/5 x r (to the 11 power), I cant find the pattern for this sequence,

OpenStudy (amistre64):

-18/5 6/5 -2/5 2/15 you sure this is right?

OpenStudy (anonymous):

yes, the answer is 2/98415, but i dont have a clue to what r is

OpenStudy (anonymous):

Try: \[-\frac{1}{3}\]

OpenStudy (amistre64):

-18 6 ------ = --- 5(-3) 5 6 -2 ---- = ---- 5(-3) 5 -2 2 ---- not= --- 5(-3) 5

OpenStudy (amistre64):

opps; lost track of the sequence ;; -1/3 is it

OpenStudy (anonymous):

yes thanks abunch

OpenStudy (amistre64):

if you divide: a{n} ----- and get a constant result; thats the answer a{n-1}

OpenStudy (anonymous):

thanks